The Auditing and Attestation (AUD) part of the CPA exam has 72 multiple choice questions with 6 options each (a through f). Let G represents the correct guesses in the test, knowing that the accountant did not study well for the test.What is the P(G=25)? What is P(G≥38)? P(16<G≤35)? P(G≤30)? Create a Probability distribution for this problem.

Please I need this question explained how to do it in Excel
Answer #1

G follows Binomial distribution with parameter n = 72 and p = 1/6 (Because there is only one option is correct out of 6 options)

Use following function in excel to find required probabilities.

1. P(G = 25) =BINOMDIST(25,72,1/6,FALSE) = 0.000101949

2. P(G≥38) = 1- P(G < 37) =1-BINOMDIST(37,72,1/6,TRUE) = 2.61957E-12

3. P(16<G≤35) = P(G≤35) - P(G≤16) =BINOMDIST(35,72,1/6,TRUE)-BINOMDIST(16,72,1/6,TRUE) = 0.08158939